OpenStudy (anonymous):

using elimination and substitution find the unknown values 2x+6y=-2 and 4x-2y=24

OpenStudy (anonymous):

do you know how to start

OpenStudy (anonymous):

no

Parth (parthkohli):

Write the first equation as \(\rm -4x -12y = 12\) then add both equations. (Elimination)

Parth (parthkohli):

For substitution, write the second equation as \(\rm 2x - y = 12\) or \(\rm y = 12 - 2x \).

OpenStudy (anonymous):

thanks :)